1. What conditions must be satisfied by b1, b2, b3, b4, and b5 for the over determined linear system to be consistent?

x1 - 4x2 = b1
x1 - 3x2 = b2
x1 + x2 = b3
x1 - 5x2 = b4
x1 + 6x2 = b5

2. Find the rank and nullity of the matrix; then verify that the values obtained satisfy Formula (4) in the Dimension Theorem.

A = 3 -2 -5 7 -29
-2 0 2 -6 14
4 7 3 19 0

a. Rank (A) = ?
b. nullity (A) = ?
c. rank (A) + nullity (A) = ?

3. For purposes of this problem, let us define a "checkerboard matrix" to be a square matrix Upper A equals left-bracket a Subscript i j Baseline right-bracket such that

a Subscript i j Baseline equals Start Layout left-brace 1st Row 1st Column 1 if i plus j is even 2nd Row 1st Column 0 if i plus j is odd End Layout

Find the rank and nullity of the following checkerboard matrix.

The 9 times 9 checkerboard matrix.

a. rank (A) =
b. nullity (A) =
1
Expert's answer
2020-01-12T15:13:20-0500

1.


"\begin{pmatrix}
 1 & -4& \ |& b_1 \\
 1 & -3 & \ |& b_2 \\
 1 & 1 & \ |& b_3 \\
 1 & -5 & \ |& b_4 \\
 1 & 6 & \ |& b_5
\end{pmatrix}"

"R_2=R_2-R_1"


"\begin{pmatrix}
 1 & -4& \ |& b_1 \\
 0 & 1 & \ |& -b_1+b_2 \\
 1 & 1 & \ |& b_3 \\
 1 & -5 & \ |& b_4 \\
 1 & 6 & \ |& b_5
\end{pmatrix}"

"R_3=R_3-R_1"


"\begin{pmatrix}
 1 & -4& \ |& b_1 \\
 0 & 1 & \ |& -b_1+b_2 \\
 0 & 5 & \ |& -b_1+b_3 \\
 1 & -5 & \ |& b_4 \\
 1 & 6 & \ |& b_5
\end{pmatrix}"

"R_4=R_4-R_1"


"\begin{pmatrix}
 1 & -4& \ |& b_1 \\
 0 & 1 & \ |& -b_1+b_2 \\
 0 & 5 & \ |& -b_1+b_3 \\
 0 & -1& \ |& -b_1+b_4 \\
 1 & 6 & \ |& b_5
\end{pmatrix}"

"R_5=R_5-R_1"


"\begin{pmatrix}
 1 & -4& \ |& b_1 \\
 0 & 1 & \ |& -b_1+b_2 \\
 0 & 5 & \ |& -b_1+b_3 \\
 0 & -1& \ |& -b_1+b_4 \\
 0 & 10 & \ |& -b_1+b_5
\end{pmatrix}"

"R_1=R_1+(4)R_2"


"\begin{pmatrix}
 1 & 0 & \ |& -3b_1+4b_2 \\
 0 & 1 & \ |& -b_1+b_2 \\
 0 & 5 & \ |& -b_1+b_3 \\
 0 & -1& \ |& -b_1+b_4 \\
 0 & 10 & \ |& -b_1+b_5
\end{pmatrix}"

"R_3=R_3-(5)R_2"


"\begin{pmatrix}
 1 & 0 & \ |& -3b_1+4b_2 \\
 0 & 1 & \ |& -b_1+b_2 \\
 0 & 0 & \ |& 4b_1-5b_2+b_3 \\
 0 & -1& \ |& -b_1+b_4 \\
 0 & 10 & \ |& -b_1+b_5
\end{pmatrix}"

"R_4=R_4+R_2"


"\begin{pmatrix}
 1 & 0 & \ |& -3b_1+4b_2 \\
 0 & 1 & \ |& -b_1+b_2 \\
 0 & 0 & \ |& 4b_1-5b_2+b_3 \\
 0 & 0 & \ |& -2b_1+b_2+b_4 \\
 0 & 10 & \ |& -b_1+b_5
\end{pmatrix}"

"R_5=R_5-(10)R_2"


"\begin{pmatrix}
 1 & 0 & \ |& -3b_1+4b_2 \\
 0 & 1 & \ |& -b_1+b_2 \\
 0 & 0 & \ |& 4b_1-5b_2+b_3 \\
 0 & 0 & \ |& -2b_1+b_2+b_4 \\
 0 & 0 & \ |& 9b_1-10b_2+b_5
\end{pmatrix}"

"4b_1-5b_2+b_3=0""-2b_1+b_2+b_4=0""9b_1-10b_2+b_5=0"

Let "b_4=r, b_5=s." Then


"4b_1-5b_2+b_3=0""-2b_1+b_2+r=0""9b_1-10b_2+s=0"

"b_3=-4b_1+5b_2""b_2=2b_1-r""-11b_1+10r+s=0"

"b_1={10 \over 11}r+{1 \over 11}s ,""b_2={9 \over 11}r+{2 \over 11}s ,""b_3={5 \over 11}r+{6 \over 11}s ,""b_4=r ,""b_5=s ,"

"r,s\in \Bbb{R}"


2.


"A=\begin{pmatrix}
 3 & -2 & -5 & 7 & -29 \\
 -2 & 0 & 2 & -6 & 14 \\
 4 & 7 & 3 & 19 & 0 \\
\end{pmatrix}"

a)

"R_1=R_1\/3"


"\begin{pmatrix}
 1 & -2/3 & -5/3 & 7/3 & -29/3 \\
 -2 & 0 & 2 & -6 & 14 \\
 4 & 7 & 3 & 19 & 0 \\
\end{pmatrix}"

"R_2=R_2+(2)R_1"


"\begin{pmatrix}
 1 & -2/3 & -5/3 & 7/3 & -29/3 \\
 0 & -4/3 & -4/3 & -4/3 & -16/3 \\
 4 & 7 & 3 & 19 & 0 \\
\end{pmatrix}"


"R_3=R_3-(4)R_1"


"\begin{pmatrix}
 1 & -2/3 & -5/3 & 7/3 & -29/3 \\
 0 & -4/3 & -4/3 & -4/3 & -16/3 \\
 0 & 29/3 & 29/3 & 29/3 & 116/3 \\
\end{pmatrix}"

"R_2=(-{3 \over 4})R_2"


"\begin{pmatrix}
 1 & -2/3 & -5/3 & 7/3 & -29/3 \\
 0 & 1 & 1 & 1 & 4 \\
 0 & 29/3 & 29/3 & 29/3 & 116/3 \\
\end{pmatrix}"

"R_1=R_1+({2 \over 3})R_2"


"\begin{pmatrix}
 1 & 0 & -1 & 3 & -7 \\
 0 & 1 & 1 & 1 & 4 \\
 0 & 29/3 & 29/3 & 29/3 & 116/3 \\
\end{pmatrix}"

"R_3=R_3-({29 \over 3})R_2"


"\begin{pmatrix}
 1 & 0 & -1 & 3 & -7 \\
 0 & 1 & 1 & 1 & 4 \\
 0 & 0 & 0 & 0 & 0 \\
\end{pmatrix}"

The rank of a matrix is the number of nonzero rows in the reduced matrix, so the rank is 2.


"rank(A)=2"

b)

Solve the matrix equation


"\begin{pmatrix}
 1 & 0 & -1 & 3 & -7 \\
 0 & 1 & 1 & 1 & 4 \\
 0 & 0 & 0 & 0 & 0 \\
\end{pmatrix}
\begin{pmatrix}
 x_1 \\
 x_2 \\
 x_3 \\
 x_4 \\
 x_5 
\end{pmatrix}=
\begin{pmatrix}
 0 \\
 0 \\
 0 
\end{pmatrix}"

If we take "x_3=t, x_4=s, x_5=u," then "x_1=t-3s+7u, x_2=-t-s-4u,"

"x_3=t,x_4=s,x_5=u."

Therefore


"X=\begin{pmatrix}
 t-3s+7u \\
 -t-s-4u \\
 t \\
 s \\
 u
\end{pmatrix}=""=
\begin{pmatrix}
 1 \\
 -1 \\
 1 \\
 0 \\
 0 
\end{pmatrix}t+
\begin{pmatrix}
 -3 \\
 -1 \\
 0 \\
 1 \\
 0
\end{pmatrix}s+
\begin{pmatrix}
 7 \\
 -4 \\
 0 \\
 0 \\
 1 
\end{pmatrix}u"

This is a null space.

Thus the nullity of the matrix A is 3.


"nullity(A)=3"

c.


"rank(A)+nullity(A)=2+3=5=n"

"A" is a matrix with "n=5" columns.


3.


"A=\begin{pmatrix}
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 \\
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1
\end{pmatrix}"

"R_3=R_3-R_1"


"\begin{pmatrix}
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 \\
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1
\end{pmatrix}"

"R_5=R_5-R_1"


"\begin{pmatrix}
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 \\
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1
\end{pmatrix}"

"R_7=R_7-R_1"


"\begin{pmatrix}
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 \\
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1
\end{pmatrix}"

"R_9=R_9-R_1"


"\begin{pmatrix}
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0
\end{pmatrix}"

"R_4=R_4-R_2"


"\begin{pmatrix}
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0
\end{pmatrix}"

"R_6=R_6-R_2"


"\begin{pmatrix}
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0
\end{pmatrix}"

"R_8=R_8-R_2"


"\begin{pmatrix}
 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 \\
 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 &0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\
 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0
\end{pmatrix}"


a.

The rank of a matrix is the number of nonzero rows in the reduced matrix, so the rank is 2.

"rank(A)=2"


b.


"rank(A)+nullity(A)=n"

"n=9"


"nullity(A)=9-2=7"

"nullity(A)=7"